# Rates of Convergence of Generalised Variational Inference Posteriors under Prior Misspecification

### v1

@@ 2.4 Assumptions
As we will be working with the infima of the different loss functions throughout, we follow [31] by adopting further notational abbreviations stemming from large deviation theory.
$L_{n}(A)\equiv\inf_{\theta\in A}\,L(X_{1}^{n},\,\theta)\,,\qquad L(A)\equiv\inf_{\theta\in A}\,L(\theta)$ (4)
As we have not fixed a metric, we follow [7, 31, 28] and use the following set as a measure of closeness:
$N_{\varepsilon}:=\{\theta\in\Theta:L(\theta)\leq L(\Theta)+\varepsilon\}\,.$ (5)
3 Main Results
We first provide the statements of the main theoretical results. We begin by characterising the GVI posteriors under bounded divergences through the empirical loss minimiser, find sufficient conditions for existence and uniqueness in infinite dimensional hypothesis spaces, and prove asymptotic consistency, culminating in a rates of convergence result. Finally, we briefly touch upon GVI posteriors under unbounded divergences, dispensing of Assumption 3. The proofs are provided in the corresponding Sections 5.1, 5.2, 5.3, 5.4, 5.5 and 5.6.
3.1 Characterising the GVI Posterior through the Empirical Loss Minimiser
Any prior within the set $\mathcal{G}\subset\mathcal{P}(\Theta)$ cannot have infinite influence under Assumptions 3 and 4, and therefore GVI prefers measures that are close to the empirical loss minimiser, such that the prior has limited influence if it does not agree with the data.
@@ 3.1 Characterising the GVI Posterior through the Empirical Loss Minimiser
Under Assumptions 1, 2, 3 and 4, for all priors $\Pi\in\mathcal{G}$, the respective GVI posterior $Q_{n}$ of Equation 3 is in the set
Moreover, suppose that $\mathcal{Q}$ is closed with respect to weak convergence and that the loss satisfies Assumption 5, then $\mathcal{R}^{\star}_{n}$ is compact with respect to weak convergence.
Significantly, this result is independent of the choice of prior from the set $\mathcal{G}$. It implies that the GVI posterior may not deviate too far from the empirical loss minimising measure, $P_{n}^{\star}$, albeit in a very rough ‘neighbourhood’. This result implies that the divergence term may never skew posteriors away from the data, given such bounded divergences, even under severe prior misspecification where the traditional Kullback–Leibler divergence were infinite. It is not even necessary for the support of the prior to intersect a subset of $\Theta$ on which the GVI posterior has full measure, nor the family $\mathcal{Q}$ and the prior $\Pi$ to be dominated by some common reference measure.
The weak compactness result allows us to characterise the existence and uniqueness of GVI posteriors. If $\Theta$ were compact then the weak compactness of $\mathcal{R}_{n}^{\star}$ above would be immediate, as any set of measures on $\Theta$ will be tight. However, as $\Theta$ will rarely be compact, we further require that the loss is coercive in order to achieve the same result.
3.2 Existence and Uniqueness of a Generalised Variational Inference Posterior
Having established that $\mathcal{R}_{n}^{\star}$ is compact with respect to weak convergence of probability measures, we are able to demonstrate the existence of minimisers of the GVI functional, Equation 3. We prove existence and uniqueness of GVI posteriors over arbitrary Polish spaces as soon as there is a single element in $\mathcal{Q}$ for which Equation 3 is finite, extending the previous results of Knoblauch [23] and Wild et al. [35]. In particular, we allow for $\Theta$ to be an infinite dimensional Polish space, rather than a finite dimensional, complete, and separable normed space. Far less do we require for $\mathcal{Q}$ to be convex; this is only required to establish uniqueness.
@@ 3.2 Existence and Uniqueness of a Generalised Variational Inference Posterior
Make Assumptions 1, 2, 3, 4, 5 and 6 and suppose also that $\mathcal{Q}$ is closed with respect to weak convergence. Then the GVI minimiser $Q_{n}$, of Equation 3, exists.
Establishing the existence of GVI posterior measures provides that the GVI objective is in fact well specified and a reasonable framework for decision makers in the nonparametric setting. While the above theorem relies on the fact that $\mathcal{R}_{n}^{\star}$ is weak compact, and hence on the boundedness of the divergence, we later establish that this assumption is not necessary given that the prior is fixed. We provide this result in Section 3.6 in the form of Proposition 8.
Unfortunately, as the set $\mathcal{Q}$ is not convex in typical situations of interest, we generally do not have uniqueness of the GVI posterior (see [29] for non–uniqueness under Kullback–Leibler approximations). For instance, the set of Gaussian measures is not convex, nor are typical mixture models [29]. However, we can make the following observation:
Corollary 3.
A convex combination of elements from $\mathcal{R}^{\star}_{n}$ is in $\mathcal{R}^{\star}_{n}$ if and only if that convex combination is also in $\mathcal{Q}$. That is, pick $\mu,\nu\in\mathcal{R}_{n}^{\star}$ and some $a\in[0,1]$, then
$\gamma:=a\mu+(1-a)\nu\in\mathcal{R}_{n}^{\star}\iff\gamma\in\mathcal{Q}\,.$
It follows as special cases, that (a) if $\mathcal{Q}$ is convex then so is $\mathcal{R}_{n}^{\star}$, and (b) if $\mathcal{Q}$ is a set of mixtures, not necessarily convex, then if individual components are contained within $\mathcal{R}_{n}^{\star}$, then so are all permitted mixtures between these. If $\mathcal{Q}$ is indeed convex then the following, stronger result on the uniqueness of the GVI minimiser then immediately follows:
Corollary 4.
Let the assumptions be as in Theorem 2. Suppose further that the divergence is strictly convex in its first argument and that $\mathcal{Q}$ is convex. Then there exists a unique minimiser of the functional $\nu\mapsto T_{n}(\nu)$.
Therefore, under the conditions of the corollary above, the GVI minimiser will be an extreme point of not only $\mathcal{R}_{n}^{\star}$ but also of $\mathcal{Q}$, exists, and is unique.
3.3 Support of the Prior is not required
Significantly, our results are independent of the choice of prior $\Pi\in\mathcal{G}$. This implies that the support of the prior need not contain the support of the GVI posterior, hence assumptions such as those in [28] where we require the prior to place positive weight on a neighbourhood of the expected loss minimiser are not necessary. Therefore, even if the prior is severely misspecified is its influence may not skew the posterior too far away from the data.
Proposition 5.
Theorem 1 holds even if $\operatorname*{supp}(Q_{n})\cap\operatorname*{supp}(\Pi)=\emptyset$.
That is, the prior and the GVI posterior may be mutually singular. Therefore, we get a stronger result than in classical and generalised Bayesian/Variational Inference that require for the prior to have positive mass on the support of the posterior measure. Since Assumption 3 applies, this will typically only be the case when the divergence is a bounded divergence rather then when $\mathcal{Q}$ and $\mathcal{G}$ are restricted. Nevertheless, this allows us to overcome challenges, as discussed in [13, 14], where Bayesian posteriors are inconsistent. Hence leading to robustness to severe prior misspecification.
This proposition is significant as it will imply that the prior need not put positive measure on Kullback–Leibler neighbourhoods of the data generating process, $P_{0}$. It therefore allows for the consistency and rate of convergence results that follow to hold in severely prior misspecified setting, even when the prior has no measure on open sets of the hypothesis space close to the loss minimisers, in the sense of $N_{\varepsilon}$.
3.4 Asymptotics and Frequentist Consistency
Theorem 1 allows us to study the asymptotics of the GVI posterior without requiring us to explicitly know it. By Theorem 2 we additionally know that the sequence of GVI posteriors exist, so we may study their asymptotic behaviour. In key difference to previous work on (generalised) Bayesian posteriors [28, 22, 31] we do not require explicit knowledge of the GVI posterior. Furthermore, when comparing our results to the frequentist consistency result of [23], the assumptions on the loss (and the data) are far less strict, and we do not require $\Theta$ to be a finite dimensional separable normed space. While the bounded divergence is a stricter requirement this additionally allows for robustness to prior misspecification as in Definition 1. To reason about the asymptotics of the sequence of GVI posterior measures we need that these exist, so we suppose in the remaining that $\mathcal{Q}$ is closed with respect to weak convergence.
A key challenge here is that we require for the family $\mathcal{Q}$ to be rich enough, that is the measures that we desire to concentrate on exists within $\mathcal{Q}$ a priori. Naturally, this challenge is non–existent in consistency results of Bayesian posteriors as $\mathcal{Q}=\mathcal{P}(\Theta)$, we however need to be able to deal with subsets of $\mathcal{P}(\Theta)$. We solve this issue by requiring for $\mathcal{Q}$ to allow for Dirac measures; by this we mean that $\mathcal{Q}$ is closed w.r.t. weak convergence of probability measures, and that for all $n$ sufficiently large it holds $P_{0}$–a.s. that $\inf_{\nu\in\mathcal{Q}}\,J_{L_{n}}(\nu)=L_{n}(\Theta)$. Note that we defined $L_{n}(A)\equiv\inf_{\theta\in A}L_{n}(\theta)$ in Equation 4.
Theorem 6.
@@ 3.4 Asymptotics and Frequentist Consistency
Make Assumptions 1, 2, 3, 4, 5, 6 and 7. Suppose further that $\mathcal{Q}$ allows for Dirac delta measures, and that $\forall A\in\mathcal{T}$ for which $L(A)>L(\Theta)$ we almost surely have that
@@
Then $P_{0}$–a.s. for any such $A\in\mathcal{T}$
$Q_{n}A\rightarrow 0.$
Moreover, suppose there exists a unique $\theta^{\star}\in\Theta$ minimising $L(\theta)$, that is $L(\theta^{\star})=L(\Theta)$ and that $\forall\theta^{\prime}\in\Theta$ where $\theta^{\prime}\neq\theta^{\star}$ $L(\theta^{\prime})>L(\theta^{\star})$, then $P_{0}$–a.s.
$Q_{n}\overset{{w}^{\star}}{\longrightarrow}\delta_{\theta^{\star}}\,.$
@@
In the moreover part we show the weak ($w^{\star}$) convergence of $(Q_{n})$ to $\delta_{\theta^{\star}}$, which we may interpret as the expected loss minimiser of $L(\theta)$, i.e. $P^{\star}_{\infty}$. The extension to infinite dimensional Polish spaces and the weakening of assumptions make this is a considerable improvement over the previous result on weak convergence in [23].
3.5 Rate of Convergence
@@
Make Assumptions 1, 2, 3, 4, 5, 6 and 7, and suppose $\mathcal{Q}$ allows for Dirac delta measures. Pick a positive sequence $\varepsilon_{n}$ such that $\varepsilon_{n}\rightarrow 0$ and $n\varepsilon_{n}\rightarrow\infty$. If
@@
Significantly, this is the first result that establishes rates of convergence for Generalised Variational Inference posterior measures. Here we allow for sequences $\varepsilon_{n}$ that converge just slower than $n^{-1}$ which should be compared to the result of [31] who finds the same rates for Bayesian posteriors with highly correlated data using Kullback–Leibler neighbourhoods. In [16] the authors finds rates of order $n^{-1/2}$ for the Bayesian posterior using stronger definitions of $N_{\varepsilon_{n}}$ through the Hellinger metric.
Moreover, since this result builds on Theorem 1 it follows that Proposition 5 applies and that the above theorem is independent of the choice of prior $\Pi\in\mathcal{G}$. This implies that we do not require $\Pi(N_{\varepsilon_{n}})>0$ for any $n\in\mathbb{N}$, and the GVI posterior may hence concentrate on hypotheses that minimise the converged loss, $L$, rather than being restricted to the support of the prior which may be misspecified.
3.6 Extensions
Throughout, we have assumed that Assumption 3 holds, that is the divergence between all possible posteriors, $\mathcal{Q}$, and all priors, $\mathcal{G}$, is bounded. This has allowed us to make significant progress in establishing asymptotic consistency and rates of convergence of GVI posterior measures; under unbounded divergences this is an open problem.
@@ 3.6 Extensions
Theorem 1 may now be satisfied with $\mathcal{Q}|_{M}$ and $\mathcal{G}=\{\Pi\}$. This, however, does not guarantee robustness to prior misspecification in the same generality as using a bounded divergence, due to the set $\mathcal{G}$ from Definition 1 being smaller. In fact, $\mathcal{Q}|_{M}$ implicitly makes assumptions on the absolute continuity of its members to the prior.
Here, we explore to what extend we may extend our results to unbounded divergences. In particular, we find that we may establish existence and uniqueness without Assumption 3 through additional work, and we make some progress on asymptotic consistency by replacing Assumption 3 by different assumptions on the divergence. Nevertheless, rates of convergence of GVI posterior measures under unbounded divergences remain an open problem.
3.6.1 Existence and Uniqueness under Unbounded Divergences
As long as there is a single element of $\mathcal{Q}$ for which the GVI functional is finite, then the posterior exists. If this further holds for all priors within the subspace $\mathcal{G}$, then it exists regardless of the prior measure.
Proposition 8.
Make Assumptions 2, 1, 6 and 5 and suppose that $\mathcal{Q}$ is closed with respect to weak convergence. Then the GVI posterior minimising Equation 3 exists.
Moreover, suppose also that the divergence is strictly convex in its first argument and that $\mathcal{Q}$ is convex. Then the GVI posterior exists and is unique.
We have now established sufficient conditions for existence and uniqueness of Generalised Variational Inference posteriors under unbounded divergences over arbitrary Polish spaces. Significantly, a GVI posterior will exist as soon as a single measure within $\mathcal{Q}$ has finite GVI objective.
3.6.2 Relaxation of $M$ and $\beta$
There is no objective reason why the different inputs into the GVI objective should be fixed with respect to incoming observations. In fact, Theorems 6 and 7 remain valid even if $\beta$, $M$, $\Pi$, or $D$ vary with $n$. We simply require that the respective assumptions are satisfied for each $n$, but these may affect the rate of convergence.
Proposition 9.
Let $\beta^{-1}:\mathbb{N}\rightarrow\mathbb{R}_{+}$ and $M:\mathbb{N}\rightarrow\mathbb{R}_{+}$ be functions on $n$ mapping to strictly positive numbers. If
$M(n)\beta^{-1}(n)n^{-1}\rightarrow 0,\quad n\rightarrow\infty$
then the conclusion of Theorem 6 is unchanged. Further, Theorem 7 holds with rates
$\varepsilon_{n}>0,\qquad\varepsilon_{n}\rightarrow 0,\qquad n\varepsilon_{n}M(n)^{-1}\beta(n)\rightarrow\infty.$
Remark.
This for instance allows for the following divergence:
$D_{TV}^{(\sqrt{n})}(Q:\Pi):=n^{1/2}||Q-\Pi||_{TV},$
preventing over–concentration around the data in cases where the prior is not severely misspecified while maintaining the results on rates of convergence. We may hence achieve an interplay between Bayesian and frequentist inference.
3.6.3 Concentration of GVI Posteriors under Arbitrary Divergences

### v2

@@ 2.2. Assumptions
As we will be working with the infima of the different loss functions throughout, we follow [shalizi2009] by adopting further notational abbreviations stemming from large deviation theory.
(2) $L_{n}(A)\equiv\inf_{\theta\in A}\,L(X_{1}^{n},\,\theta)\,,\qquad L(A)\equiv\inf_{\theta\in A}\,L(\theta)$
As we have not fixed a metric, we follow [berk1966, shalizi2009, miller2021] in using:
(3) $N_{\varepsilon}:=\{\theta\in\Theta:L(\theta)\leq L(\Theta)+\varepsilon\}\,.$
2.3. Misspecification of the Model and the Prior
As identified in [walker2013, shalizi2009] the elements of $\Theta$ that the Bayesian posterior targets are those minimising the Kullback–Leibler divergence between the data generating process and the hypothesis measure. Denoting this as a function on $\Theta$, we write:
$h(\theta)=\lim_{n\rightarrow\infty}\frac{1}{n}\mathbb{E}_{P_{0}}\left[\log\frac{p_{0}(x_{1}^{n})}{p_{\theta}(x_{1}^{n})}\right]$
which may possibly be infinite; note that the expectation is with respect to $P_{0}$ and hence over $\Omega$. This is the loss function targeted in classical Bayesian inference, in particular the elements of $\Theta$ that minimise the above minimise the Kullback–Leibler divergence between the data generating process and the hypothesis measure. However, this may be suboptimal in many cases where the model is misspecified, for instance not containing the data generating process, (see e.g. [bernardo2000, mclatchie2025]). It may therefore be desirable to instead minimise a different robust divergence, such as a density power divergence [basu1998], but we do not necessitate the loss to approximate one and it may be more general.
Moreover, if the prior has full measure on the set of elements where this KL rate is infinite, that is on the set $I:=\{\theta\in\Theta:h(\theta)=\infty\}$, then all previous results on consistency, let alone rates of convergence, of (Generalised) Bayesian inference fail [shalizi2009]. In this case we say that the prior is severely misspecified. In this we also include priors that lead to inconsistent posteriors such as those observed in [diaconis1986b]. To make this rigorous, we place the following definition of robustness to prior misspecification.
Definition 1.
Classical results on Bayesian sensitivity analysis deal with small perturbations of some reference prior by some other measure through additive or multiplicative misspecification (see [dey1994, gustafson1995]). That is some reference prior $\Pi$ is perturbed by another probability measure $G\in\mathcal{P}(\Theta)$ through a constant $\varepsilon\in[0,1]$, e.g. $\Pi_{\varepsilon}=(1-\varepsilon)\Pi+\varepsilon G$. Our results are distinct to this line of work as we deal with asymptotic inconsistency, and we will not use such explicit formulations for $\mathcal{G}$. We discuss this distinction in Appendix B.
2.4. Generalised Variational Inference
The Bayesian posterior—consisting of the likelihood function of the data, $p_{\theta}(X_{1}^{n})$, and a prior measure, $\Pi$—is the random measure
$\Pi_{n}A:=\cfrac{\int_{A}p_{\theta}(X_{1}^{n})\,\Pi(d\theta)}{\int_{\Theta}p_{\theta}(X_{1}^{n})\,\Pi(d\theta)}=\cfrac{\int_{A}R_{n}(\theta)\,\Pi(d\theta)}{\int_{\Theta}R_{n}(\theta)\,\Pi(d\theta)}$
as usual. It is immediate that we may replace $p_{\theta}(X_{1}^{n})$ by $R_{n}(\theta):=p_{\theta}(X_{1}^{n})/p_{0}(X_{1}^{n})$ without changing $\Pi_{n}$ to see why we consider $h(\theta)$ above. In particular, under mild conditions, the generalised Shannon–MacMillan–Breiman theorem [algoet1988] guarantees that $n^{-1}\log R_{n}(\theta)\rightarrow-h(\theta)$ (a.s.). If we instead replace the likelihood, $p_{\theta}(X_{1}^{n})$, by the exponential of a negative loss acting on the data, that is by $\exp\{-nL(X_{1}^{n},\theta)\}$, we get Generalised Bayesian Inference (GBI) [bissiri2016, miller2021], or a Gibbs posterior [alquier2016]. If we discard the prior, we may instead get an energy–based, frequentist model.
As shown in [zellner1988] Bayes may be equivalently formulated as an optimisation problem where $\Pi_{n}$ is the measure minimising the ‘Evidence Lower Bound’—the expectation of the negative log–likelihood, and the Kullback–Leibler divergence to the prior—over all probability measures acting on the hypothesis space. Tangentially, minimising a divergence other than the KL divergence to the Bayesian posterior may result in nicer optimisation problems in practice, this however does not nicely decompose into an expectation acting on the loss and a divergence measure acting on the prior, conflating model and prior misspecification; see also section 2.3.2 in [jeremias2022]. GVI instead replaces the negative log–likelihood by any loss function, similar to GBI, as well as the KL divergence by any divergence, generalising these previous approaches. Typical choices of loss functions are those that either offer attractive computational benefits, and those that incorporate the hypothesis measure robustly, including those based on robust divergences; see [mildner2025].
Having introduced the necessary assumptions in Section 2.2 we are able to establish that GVI posteriors over arbitrary Polish spaces minimising $\nu\mapsto T_{n}(\nu)$ exist and, under further assumptions, are also unique, extending the previous results of Knoblauch [knoblauch2021] and Wild et al. [wild2023]. In particular, we allow for $\Theta$ to be an infinite dimensional Polish space, rather than a finite dimensional, complete, and separable normed space. Far less do we require for $\mathcal{Q}$ to be convex; this is only required to establish uniqueness.
@@ 2.4. Generalised Variational Inference
Make Assumptions 1, 3, 4 and 2. Suppose also that $\mathcal{Q}$ is closed with respect to weak convergence. Then a GVI posterior minimising Equation 1 exists.
If it further holds that the divergence is strictly convex in its first argument and that $\mathcal{Q}$ is convex, then there exists a unique minimiser to the functional $\nu\mapsto T_{n}(\nu)$.
Establishing the existence of GVI posterior measures provides that the GVI objective is in fact well specified and a reasonable framework for decision makers in the nonparametric setting.
3. Main Results
By Theorem 1 we know that the sequence of GVI posteriors exist, so we may study their asymptotic behaviour. In key difference to previous work on (generalised) Bayesian posteriors [miller2021, kleijn2006, shalizi2009] we do not require explicit knowledge of the GVI posterior. Furthermore, when comparing our results to the frequentist consistency result of Knoblauch [knoblauch2021], the assumptions on the loss (and the data) are far less strict, and we do not require $\Theta$ to be a finite dimensional separable normed space. While the bounded divergence is a stricter requirement this additionally allows for robustness to prior misspecification as in Definition 1 and, given the right choice of divergence, hold for all priors $\Pi\in\mathcal{P}(\Theta)$.
To reason about the asymptotics of the sequence of GVI posterior measures we need that these exist, so we suppose in the remaining that $\mathcal{Q}$ is closed with respect to weak convergence. A key challenge here is that we require for the family $\mathcal{Q}$ to be rich enough, that is the measures that we desire to concentrate on exist within $\mathcal{Q}$ a priori. Naturally, this challenge is non–existent in consistency results of Bayesian posteriors as $\mathcal{Q}=\mathcal{P}(\Theta)$, we however need to be able to deal with subsets of $\mathcal{P}(\Theta)$. We solve this issue by requiring for $\mathcal{Q}$ to allow for Dirac measures; by this we mean that $\mathcal{Q}$ is closed with respect to the weak convergence of probability measures, and that for all $n$ sufficiently large it holds $P_{0}$–a.s. that $\inf_{\nu\in\mathcal{Q}}\,J_{L_{n}}(\nu)=L_{n}(\Theta)$. Recall that we defined $L_{n}(A)\equiv\inf_{\theta\in A}L_{n}(\theta)$ in Equation 2.
@@ 3. Main Results
Make Assumptions 1, 2, 5, 3, 4 and 6. Suppose further that $\mathcal{Q}$ allows for Dirac delta measures, and that $\forall A\in\mathcal{T}$ for which $L(A)>L(\Theta)$ we almost surely have that
@@
Then for any such $A\in\mathcal{T}$ it $P_{0}$–a.s. holds that $Q_{n}A\rightarrow 0$.
Moreover, suppose there exists a unique $\theta^{\star}\in\Theta$ minimising $L(\theta)$, then
$Q_{n}\overset{{w}^{\star}}{\longrightarrow}\delta_{\theta^{\star}}\,$
$P_{0}$–a.s.
@@
In the moreover part we show the weak ($w^{\star}$) convergence of $(Q_{n})$ to $\delta_{\theta^{\star}}$, which we may interpret as the expected loss minimiser of $L(\theta)$, i.e. $P^{\star}_{\infty}$. The extension to infinite dimensional Polish spaces and the weakening of assumptions make this is a considerable improvement over the previous result on weak convergence in [knoblauch2021].
@@
Make Assumptions 1, 2, 5, 3, 4 and 6, and suppose $\mathcal{Q}$ allows for Dirac delta measures. Pick a positive sequence $\varepsilon_{n}$ such that $\varepsilon_{n}\rightarrow 0$ and $n\varepsilon_{n}\rightarrow\infty$. If
@@
Significantly, this is the first result that establishes rates of convergence for Generalised Variational Inference posterior measures. Here we allow for sequences $\varepsilon_{n}$ that converge just slower than $n^{-1}$ which should be compared to the result of [shalizi2009] who finds the same rates for Bayesian posteriors with highly correlated data using Kullback–Leibler neighbourhoods. In [ghosal2007] the authors finds rates of order $n^{-1/2}$ for the Bayesian posterior using stronger definitions of $N_{\varepsilon_{n}}$ through the Hellinger metric.
Moreover, given that Assumption 5 applies, we find that the above theorems are independent of the choice of prior $\Pi\in\mathcal{G}$. This implies that we do not require $\Pi(N_{\varepsilon_{n}})>0$ for any $n\in\mathbb{N}$, given $\mathcal{G}=\mathcal{P}(\Theta)$, and the GVI posterior may hence concentrate on hypotheses that minimise the converged loss, $L$, rather than being restricted to the support of the prior which may be misspecified.
There is no objective reason why the different inputs into the GVI objective should be fixed with respect to incoming observations. In fact, Theorems 2 and 3 remain valid even if $\beta$, $M$, $\Pi$, $L$, or $D$ vary with $n$. We simply require that the respective assumptions are satisfied for each $n$, but these may affect the rate of convergence.
Theorem 4.
Theorems 2 and 3 remain valid if $\Pi$, $\Theta$, $\mathcal{P}$}, $\mathcal{Q}$, $L(\cdot,\cdot)$, $M$, $D(\cdot:\cdot)$, and $\beta$ depend on $n$ but satisfy the respective assumptions for each $n$. In particular, we require $M_{n}\beta_{n}^{-1}n^{-1}\rightarrow 0$, and that the sequence $\varepsilon_{n}$ satisfies $n\varepsilon_{n}M_{n}^{-1}\beta_{n}\rightarrow\infty$.
Moreover, Theorems 2 and 3 hold regardless of the prior $\Pi\in\mathcal{G}$, and we may pick the divergence $D$ such that $\mathcal{G}=\mathcal{P}(\Theta)$.
Remark.
Notably, this allows for situations in which $Q_{n}$ and $\Pi$ are mutually singular, and hence whenever the prior is severely misspecified.
As another extension of the preceding theorems, it is certainly permissible for the prior to put mass outside the space $\Theta$. In fact, we may even allow for its entire mass to be outside of $\Theta$, in which case the divergence is maximal. Such situations may arise when the space $\Theta$ is embedded inside some larger space $T$. Consider the case where $\Theta$ corresponds to some model class $\mathcal{P}\subset\mathcal{P}(\mathbb{R})$ and the prior is in fact over the entire space $\mathcal{P}(\mathbb{R})$.
4. Finite Sample Bound and Extensions
4.1. A Finite Sample Bound
Before we bound the objective of the GVI posterior under arbitrary priors $\Pi\in\mathcal{G}$, we define an empirical loss minimising measure corresponding to a minimiser of the functional $\nu\mapsto J_{L_{n}}(\nu)$.
Definition 2.
We denote $P_{n}^{\star}\in\mathcal{Q}$ such that $\forall Q\in\mathcal{Q}$ it holds that $J_{L_{n}}(P_{n}^{\star})\leq J_{L_{n}}(Q)$. This is guaranteed to exist under Assumptions 1, 2 and 3 supposing that $\mathcal{Q}$} is weak closed.
Proposition 5.
Under Assumptions 1, 2, 5 and 3, for all priors $\Pi\in\mathcal{G}$, the respective GVI posteriors $Q_{n}$ of Equation 1 are in the set
Moreover, suppose that $\mathcal{Q}$ is closed with respect to weak convergence, then $\mathcal{R}^{\star}_{n}$ is compact with respect to weak convergence.
This result provides an instrumental role in proving the theorems of the preceding section.
@@ 4.1. A Finite Sample Bound
Proposition 5 may now be satisfied with $\mathcal{Q}|_{M}$ and $\mathcal{G}=\{\Pi\}$. This, however, does not guarantee robustness to prior misspecification in the same generality as using a bounded divergence, due to the set $\mathcal{G}$ from Definition 1 being smaller. In fact, $\mathcal{Q}|_{M}$ implicitly makes assumptions on the absolute continuity of its members to the prior.
4.2. Generalisation Performance
Let $P^{n}_{0}\equiv P_{0}(X_{n}|\sigma(X_{1}^{n-1}))$ be the conditional under the data generating process, and denote the posterior predictive $Q^{n}_{P_{\theta}}\equiv\int_{\Theta}P_{\theta}(X_{n}|\sigma(X_{1}^{n-1}))Q_{n}(d\theta)$ under the GVI posterior $Q_{n}$. It follows from [shalizi2009] that the following result on the generalisation performance of GVI holds.
Theorem 6.
Let $L(X_{1}^{n},\,\theta):=n^{-1}\log(p_{0}(X_{1}^{n})/p_{\theta}(X_{1}^{n}))$, and suppose that the assumptions of Theorem 2 hold. Then it $P_{0}$–a.s. holds that
$\limsup_{n\rightarrow\infty}d_{H}^{2}(P_{0}^{n},Q^{n}_{P_{\theta}})$ $\leq L(\Theta)$
$\limsup_{n\rightarrow\infty}d_{tv}^{2}(P_{0}^{n},Q^{n}_{P_{\theta}})$ $\leq 4L(\Theta)$
where $d_{H}$ and $d_{tv}$ are respectively the Hellinger and total variation distances.
Note that this result holds independently of the choice of prior measure $\Pi\in\mathcal{P}(\Theta)$, and $L(\Theta)$ corresponds to $h(\Theta)\equiv\inf_{\theta\in\Theta}h(\theta)$ from Section 2.3.
4.3. On Consistency of GVI Posteriors under Arbitrary Divergences